40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7149     Entry time: Fri Aug 10 19:49:11 2012
Author: Eric 
Type: Summary 
Category: Locking 
Subject: Michelson Locking Procedure and Measurements 

Today I worked on locking the Michelson. Here's what I did:

 

  1. Open Data Viewer and Restore Settings /users/Templates/JenneLockingDataviewer/MICH.xml. This opens the C1:LSC-ASDC_OUT and C1:LSC-AS55_Q_ERR plots.

  2. Check the LSC screen to verify that the path between the Servo Filter Modules and the SUS Ctrls are outlined in green. If not turn on the OUT button within the Filter Servo Modules, enable LSC mode, and turn on the SUS Ctrls for the BS.

  3. Misalign all optics other than BS and one of ITMX and ITMY. The ITMY was already well-aligned from my work on locking the YARM, so I actually chose to misalign ITMY at first.

  4. Restore BS and ITMX. Use the AS camera on the video screen as your guide when aligning ITMX.

  5. Adjust pitch and yaw of ITMX until a bright, circular spot appears near the middle of the AS camera.

  6. Now restore ITMY and adjust pitch and yaw until a second circular spot appears on the AS camera.

  7. Adjust both ITMX and ITMY until both bright spots occupy the same location. If the spots remain bright when they are in the same location you are locking onto a bright fringe actually, and need to flip the sign of the gain on the MICH servo filter modules. I had to do this today in fact, as discussed in ELOG 7145.

  8. If the sign is correct, the two beams should interfere destructively and the formerly bright spots will form a comparatively dark spot. The shape of the spot will likely be two bright lobes separated by a dark middle.

  9. C1:LSC-ASDC_OUT should be a roughly flat signal, and the goal now is to minimize the magnitude of this signal. The smaller this signal, the darker the AS camera should look. Decent target values for C1:LSC-ASDC_OUT are around 0.10 to 0.05.

 

Once I did this, I made measurements by exciting C1:SUS-ITMY_LSC_EXC and measuring with C1:LSC-AS55_Q_ERR. I ran a logarithmic swept sine response from 1 to 1000 Hz again, with an envelope amplitude dependence. Again I looked at the measured transfer function and coherence. I was able to get good coherence, but it was somewhat erratic in that it dipped low at high frequency multiple times.

ELOG V3.1.3-